# Self-Pro: Self-Prompt and Tuning Framework for Graph Neural Networks

### v2

@@ Abstract
Graphs have become an important modeling tool for Web applications, and graph neural networks (GNNs) have achieved great success in graph representation learning. However, their performance heavily relies on a large amount of supervision. Recently, “pre-train, fine-tune” has become the paradigm to address the issues of label dependency and poor generalization. However, the pre-training strategies vary for graphs with homophily and heterophily, and the objectives for various downstream tasks also differ. This leads to a gap between pretexts and downstream tasks, resulting in “negative transfer” and poor performance. Inspired by prompt learning in natural language processing, many studies turn to bridge the gap and fully leverage the pre-trained model. However, existing methods for graph prompting are tailored to homophily, neglecting inherent heterophily on graphs. Meanwhile, most of them rely on randomly initialized prompts, which negatively impact on the stability. Therefore, we propose Self-Prompt, a prompting framework for graphs based on the model and data itself. We first introduce asymmetric graph contrastive learning as pretext to address heterophily and align the objectives of pretext and downstream tasks. Then we reuse the component from pre-training as the self adapter and introduce self-prompts based on graph itself for task adaptation. Finally, we conduct extensive experiments on 11 benchmark datasets to demonstrate its superiority. We provide our codes at https://github.com/gongchenghua/Self-Pro.
@@ 1 Introduction
The complex relations in real word can be modeled using graphs. The powerful modeling capability of graphs enables the realization of web applications such as community fraud [11] and automated account detection [32] or drug development [2] and discovery of biological structures [37]. Meanwhile, GNNs have achieved great success in representation learning of graph-structured data [12, 28, 39, 16]. However, one of the fundamental problems for traditional GNNs is that they heavily depend on task-specific labels as supervision, which are costly and difficult to obtain. To solve the aforementioned problems, many studies turn to “pre-train, fine-tune” strategies [8]. Through self-superviesed pretexts, GNNs are pre-trained without labels and then fine-tuned with labeled data. However, the “pre-train, fine-tune” paradigm empirically faces the following challenges:
$\bullet$ (C1) Existing pre-training methods on graph explicitly or implicitly follow the homophily assumption [17], which further restricts the generalization of GNNs on graphs with more diverse properties (e.g., heterophily). So how to find a more generalized pre-training method accommodating more diverse graphs, such as graphs with homophily and heterophily?
$\bullet$ (C2) The objective gap between pre-training and fune-tuning phase may result in “negative transfer” [30] problem. For example, typical pretext on graphs utilize link prediction [13] to map the representation of adjacent nodes closer in latent space. However, downstream node classification focuses on the similarity of the intra-class and inter-class cluster. So how to reformulate pretexts and downstream tasks into the same template to avoid the negative knowledge transfer?
@@
Inspired by the success of prompt learning in NLP, “pre-train, prompt” has been introduced to the graph domain. The naive language prompt in NLP refers to transforming the downstream tasks into text generation template by appending prompt text since pre-trained language models have powerful text generation capabilities. For example, a sentiment analysis task for the statement “The ECML-PKDD 2024 will accept many interesting papers!” can be transferred to a text generation task by appending prompt like “The ECML-PKDD 2024 will accept many interesting papers! That is so [·] .” The language model is prone to make the prediction like “exciting” rather than “bad” in the blank. Through alignment of objectives between upstream and downstream, the pretrained models can be fully leveraged and perform well even in few-shot scenarios.
Although some works [20, 24, 25, 27, 3] have attempted to design prompt framework for graphs, most of them introduce randomly initialized parameters as virtual prompt. This greatly limits the generalization and stability of prompt learning on graphs and poses a new challenge:
$\bullet$ (C4) Existing graph prompt methods mostly rely on randomly initialized virtual prompt, which neglects the rich information inherently contained in the graph. So how to design prompt based on graph itself to guide the model to better leverage the pre-training knowledge?
@@
$\bullet$ We propose a self-prompt strategy for Self-Pro, hinging on the self adapter, the semantic and structural prompts to provide more prompt information and transfer the pre-trained knowledge for downstream adaptation.
@@ 2.1 Graph Neural Networks
Recently, GNNs have received significant attention for web applications and many advanced GNNs are proposed [12, 28] on homophilic graphs. Most of them follow the message-passing framework, where nodes update their representations by combining and aggregating messages from their neighbors. To cope with more complex patterns, many studies have designed GNNs for heterophilious graphs [4, 16, 22, 39]. The key of GNNs for heterophily is to discriminate heterophilic neighbors and utilize multi-hop neighbors. However, these GNNs all follow the supervised learning paradigm and heavily rely on labels. To tackle the issue, many studies have shifted their focus towards unsupervised graph learning and the “pre-train, fine-tune” paradigm [33].
@@ 2.2 Graph Pre-training
Having witnessed the success of pre-training in language and vision fields, a myriad of graph pre-training methods have emerged. Graph pre-training models mine intrinsic properties on graphs through self-supervised pretexts, and then are fine-tuned for downstream tasks. Graph pre-training strategies [33] can be broadly categorized into types: generative and contrastive methods. Generation-based methods including various Graph Auto-Encoders (GAEs) [13, 7, 14, 15] use the input as the supervision signal and reconstruct the feature and structure of graph. Contrast-based methods including various Graph Contrastive Learning (GCL) frameworks [29, 40, 6, 35] construct different views for graphs through data augmentation or prior knowledge and then maximize their agreement. It is worth noting that most graph pre-training methods are exclusively applicable to graphs with homophily, only a few works [19, 35, 34] have taken into account the extensive heterophily on graphs. Meanwhile, there is an inherent gap between the objectives of pre-training and fine-tuning phase. Pretexts extract general knowledge from graphs without supervision, while downstream tasks are rooted in specific labels. This gap may have a negative impact on knowledge transfer [30] and hurt the performance of downstream tasks.
@@ 2.3 Graph Prompt Learning
Originated from NLP, “pre-train, prompt” paradigm [18] reformulates downstream tasks into the same template of pretexts and designs prompts for downstream adaptation. Prompt learning fully unleashes the potential of pre-trained models where adjusting only a few parameters can achieve great results with limited labels. Inspired by the success of prompt learning in nature language [1, 31] and visual data [9, 10], the graph domain has shifted the focus towards “pre-train, prompt” paradigm [26]. GPPT [24] and GraphPrompt [20] both align the objective of downstream tasks with link prediction. ProG [25] reformulate different-level tasks to graph level and further introduce virtual prompt graphs with meta-learning. GPF [3] injects learnable prompts into the feature space for task adaptation while VNT-GPPE [27] adjusts pre-trained graph transformers by prompting virtual nodes. Similar to graph pre-training, existing graph prompting methods only focus on homophilic graphs while overlook heterophilic graphs, which restricts their wide applications in the real word. Meanwhile, they heavily rely on randomly initialized prompts, and the added virtual features or nodes are difficult to interpret. All of them ignore the rich information inherent in graph itself and the potential of prompt within models.
@@ 3 Preliminaries
Pre-train, Fine-tune.
The goal of graph pre-training is to learn a GNN encoder $f_{\theta}$ parameterized by $\theta$ to map each node $v$ to a low-dimensional representation $\mathbf{h}_{v}$, then $\mathbf{h}_{v}$ are fine-tuned according to the requirements of downstream tasks.
Graph Contrastive Leaning (GCL).
GCL with smoothing [36] is one GCL schema which ensures that neighboring nodes the have similar representations, the objective of pretext of is:
@@
where $\mathbf{h}_{v},\mathbf{h}_{v^{+}},\mathbf{h}_{v^{-}}$ are representations of nodes $v,v^{+}$ and $v^{-}$ obatined by $f_{\theta}(G)$, and $\tau$ is the temperature hyper-parameter. $\mathcal{N}(v)$ is the positive sample set containing one-hop neighborhoods of node $v$ and $\mathcal{V^{-}}$ is the negative sample set randomly sampled. The illustration for this scheme of smoothing is presented in Fig. 1(a). Another GCL schema [40, 35] shown in Fig. 1(b) learns representations by contrasting graphs under augmented views. Specifically, the representation of node $v$ in one augmented view is learned to be close to that from the other view and far away from the representations of other nodes. Given two augmented $G_{1}$ and $G_{2}$ from $G$, the objective of pretext is as follows:
@@ 4.1 Pre-training Method
For graph pre-training, contrastive methods provide broader applicability and overlapping task sub-spaces for better knowledge transfer [21]. However, most GCL schemes explicitly or implicitly follow the homophily assumption [17, 5] and empirically face issues on heterophilic graphs. So we choose the Graph Asymmetric Contrastive Learning (GraphACL) [34] as our backbone to accommodate graph with heterophily and homophily. As shown in Fig. 1(c), GraphACL introduces the asymmetric framework with a projector. Unlike traditional symmetric schemes, the design of projector allows neighboring nodes to have different representations, and it has been theoretically and experimentally proven in [34] to capture high-order neighbor similarities on graphs and effectively handle heterophily. The pretext loss of GraphACL is as follow:
@@
where $\mathbf{z}_{v}=g(\mathbf{h}_{v})$ and $g(\cdot)$ can be a linear transformation layer or a MLP. The positive sample set is from $\mathcal{N}(v)$ and random data augmentation are not required. $\mathbf{h}_{v}$ and $\mathbf{h}_{v^{+}}$ are obtained separately from two encoders $f_{\theta}$ and $f_{\delta}$ updated by exponential moving average to model more complex patterns and prevent $g(\cdot)$ from degenerating to the identity function [34]. The projector enables the model to capture higher-order context information of nodes. Similar to GCL with augmented views, GraphACL also abandons the projector and solely uses the output of encoder for downstream tasks. However, the projector also contains rich pre-trained knowledge, simply discarding it would be inappropriate. It can be combined with prompt techniques to empower downstream tasks.
@@ 4.2 Prompt Template
Prompt learning in NLP reformulates both upstream and various downstream tasks into the unified template of masked language modeling [1, 18]. As for GCL, the core is to maximize the mutual information between positive sample pairs and minimize that between negative sample pairs. To align the pretext with downstream tasks, we focus on the most common downstream tasks on graphs: link prediction and node classification, and propose a unified prompt template, denoted as $p(\cdot)$. The template transforms various tasks into pairwise similarity calculation between tokens:
@@
In the pre-training phase, the targets are set as positive and negative samples for anchor nodes. We use dot product in $p(\cdot)$ and the objective is to minimize the distance between positive sampled pairs and maximize that between negative sampled pairs in the representation space as illustrated in Eq. 3.
@@
For link prediction, the targets are set as adjacent and non-adjacent nodes for anchor nodes. The objective is to minimize the distance between pairs of adjacent nodes and maximize that between pairs of non-adjacent nodes in the representation space.
@@
Unlike traditional node classification, we set prototype for each class rather than predict solely by direct labels. Through class prototypes as targets, the node classification can be reformulated into pairwise similarity calculation. The objective is to minimize the distance between nodes and intra-class prototype, while maximize that between nodes and inter-class prototype. Formally, we denote the set of class prototype tokens as $\mathcal{C}=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{C}\}$, where $C$ is the number of classes. Each prototype token shares the same dimension with the node token and preserves class-specific semantics. Each prototype is initialized by the mean token of labeled nodes for each class [24, 20, 21]:
@@ 4.3 Self-Prompt Architecture
Most “pre-train,fine-tune” approaches for graphs introduce an additional layer for the downstream fine-tuning. Meanwhile, the projector in GCL plays a crucial role [5] in the pretext and contains partial knowledge obtained during pre-training. For example, the projector in GraphACL [34] is to preserve the context information of nodes and model local neighborhood distribution without homophily assumption. However, the projector is always discarded after the pre-trainin phase. This motivates us to further reuse the projector as the self adapter and construct a naturally parameterized architecture for tuning. Given the input graph $\mathcal{G}$, an encoder $f_{\theta}$ parameterized by $\theta$ and the projector $g_{\phi}$ parameterized by $\phi$, the pre-training objective can be formally formulated as:
@@
where $\mathcal{L}_{dow}$ is optimized by the supervision and will be introduced in Section 4.4. Note the parameters of projector $g_{\phi}$ are initialized from pre-training stage, denoted as $g_{\phi}^{*}$. By reusing the projector as a self adapter for downstream tasks, the pre-training knowledge can be retained and smoothly transferred to the downstream, greatly improving the efficiency of tuning for specific tasks without additional parameters. More complexity analysis is presented in Appendix 0.A.
@@ 4.4 Self-Prompt Injection and Tuning
Although the objectives of pretext and downstream tasks are aligned, the task-specific prompt is still required for downstream adaptation. Currently, most works introduce trainable vectors with random initialization for prompting. The specific forms include injecting prompts in the feature space [3] and representation space [20], or using virtual nodes [27] or graphs [25] as prompts. This strategy is not stable and relies on the initialization, motivating us to design more stable prompts. Considering two key elements of the graph, namely attributes and structure, we approach the problem from these two perspectives and design prompts based on the rich information inherent in the graph itself.
@@
Instead of introducing additional parameters, the original input can be modified to provide self prompts. From structural perspective, we can add or remove edges on the adjacency matrix based on prior knowledge for prompt. There have been studies [39, 16] emphasizing the importance of multi-hop neighbors on graphs. Especially, most nodes share the same label with their two-hop neighbors on most homophilic graphs [34]. Therefore, we can simply add two-hop neighbors into the adjacency matrix as self structural prompt for node classification. Formally, we define the two-hop graph $\mathcal{G}_{2}$ from $\mathcal{G}$ and its adacency matrix is $\mathbf{A}_{2}$, where $(\mathbf{A}_{2})_{ik}=1$ if there exists $j$ such that $e_{ij}\in\mathcal{E}$ and $e_{jk}\in\mathcal{E}$. Then we feed the $\mathcal{G}_{2}$ into our frozen pre-trained encoder $f_{\theta}$ to obtain token $\mathbf{t}_{v}$ of node $v$ for tuning:
@@
Compared with the graph structure, node attributes can be even more important in certain situations, such as on heterophilic graph. But how to transform the semantic of attributes into the representation space is a challenge. Inspired by [35], we construct a unit matrix $\mathbf{I}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ to replace the adjacency matrix $\mathbf{A}$ of the original graph $\mathcal{G}$. Then we feed the prompt graph $\mathcal{G}_{I}$ into the pre-trained encoder $f_{\theta}$ to obtain semantic token $\mathbf{s}_{v}$:

### v3

# Self-Pro: A Self-Prompt and Tuning Framework for Graph Neural Networks

@@ Abstract
Graphs have become an important modeling tool for web applications, and Graph Neural Networks (GNNs) have achieved great success in graph representation learning. However, the performance of traditional GNNs heavily relies on a large amount of supervision. Recently, “pre-train, fine-tune” has become the paradigm to address the issues of label dependency and poor generalization. However, the pre-training strategies vary for graphs with homophily and heterophily, and the objectives for various downstream tasks also differ. This leads to a gap between pretexts and downstream tasks, resulting in “negative transfer” and poor performance. Inspired by prompt learning in Natural Language Processing (NLP), many studies turn to bridge the gap and fully leverage the pre-trained model. However, existing methods for graph prompting are tailored to homophily, neglecting inherent heterophily on graphs. Meanwhile, most of them rely on the randomly initialized prompts, which negatively impact on the stability. Therefore, we propose Self-Prompt, a prompting framework for graphs based on the model and data itself. We first introduce asymmetric graph contrastive learning for pretext to address heterophily and align the objectives of pretext and downstream tasks. Then we reuse the component from pre-training phase as the self adapter and introduce self-prompts based on graph itself for task adaptation. Finally, we conduct extensive experiments on 11 benchmark datasets to demonstrate its superiority. We provide our codes at https://github.com/gongchenghua/Self-Pro.
@@ 1 Introduction
The complex relations in the real world can be modeled using graphs. The powerful modeling capability of graphs enables the realization of web applications such as community fraud [11] and automated account detection [32] or drug development [2]. Meanwhile, GNNs have achieved great success in representation learning of graph-structured data [12, 28, 38, 16]. However, one of the fundamental problems for traditional GNNs is that they heavily depend on task-specific labels as supervision, which are costly and difficult to obtain. To solve the aforementioned problems, many studies turn to “pre-train, fine-tune” strategies [8]. Through self-supervised pretexts, GNNs are pre-trained without labels and then fine-tuned with labeled data. However, the “pre-train, fine-tune” paradigm empirically faces the following challenges:
$\bullet$ (C1) Existing pre-training methods for graph explicitly or implicitly follow the homophily assumption [17], which further restricts the generalization of GNNs on graphs with more diverse properties (e.g., heterophily). So how to find a more generalized pre-training method accommodating more diverse graphs, such as graphs with homophily and heterophily?
$\bullet$ (C2) The gap between objectives of the pre-training and fune-tuning phase may result in “negative transfer” [30] problem. For example, a typical pretext on graphs utilizes the link prediction [13] to map representation of adjacent nodes closer in the latent space. However, downstream task such as node classification focuses on the similarity of the intra-class and inter-class clusters. So how to reformulate the pretexts and downstream tasks into same template to avoid the negative knowledge transfer?
@@
Inspired by the success of prompt learning in NLP, “pre-train, prompt” has been introduced to the graph domain. The naive language prompt in NLP refers to transforming the downstream tasks into text generation template by appending prompt text since pre-trained language models have powerful text generation capabilities. For example, a sentiment analysis task for the statement “The ECML-PKDD 2024 will accept many interesting papers!” can be transferred to a text generation task by appending prompt like “The ECML-PKDD 2024 will accept many interesting papers! That is so [__].” The language model is prone to make the prediction like “exciting” or “wonderful” rather than “bad” or “sad” in the blank. Through alignment of objectives between pretext and downstream tasks, the pretrained models can be fully leveraged and perform well even in the few-shot scenarios. Although some works [20, 24, 25, 27, 3] have attempted to design the prompt framework for graphs, most of them introduce the randomly initialized parameters as virtual prompt. This greatly limits the generalization and stability of prompt learning on graphs and poses a new challenge:
$\bullet$ (C4) Existing graph prompt methods mostly rely on the randomly initialized virtual prompt, which neglects the rich information inherently contained in the graph. So how to design prompt based on graph itself to guide the model to better leverage the pre-trained knowledge?
@@
$\bullet$ We propose a self-prompt strategy for Self-Pro, hinging on the self adapter, semantic and structural prompts to provide extra information and transfer the pre-trained knowledge for downstream adaptation.
@@ 2.1 Graph Neural Networks
Recently, GNNs have received significant attention for real-world applications and many advanced GNNs are proposed [12, 28] on homophilous graphs. Most of them follow the message-passing framework, where nodes update their representations by combining and aggregating messages from their neighbors. To cope with more complex patterns, many studies have designed GNNs for heterophilous graphs [4, 16, 22, 38]. The core idea of GNNs for heterophily is to discriminate heterophilous neighbors and utilize multi-hop neighbors. However, these GNNs all follow the supervised learning paradigm and depend on labels heavily. To tackle this issue, many studies have shifted their focus towards unsupervised graph learning and the “pre-train, fine-tune” paradigm [33].
@@ 2.2 Graph Pre-training
Having witnessed the success of pre-training in language and vision fields, a myriad of graph pre-training methods have emerged. Graph pre-training models mine intrinsic properties on graphs through self-supervised pretexts, and then are fine-tuned for downstream tasks. Graph pre-training strategies [33] can be broadly categorized into two types: generative and contrastive methods. Generation-based methods including various Graph Auto-Encoders (GAEs) [13, 7, 14, 15] use the input as the supervision signal and reconstruct the feature and structure of graph. Contrast-based methods including various Graph Contrastive Learning (GCL) frameworks [29, 39, 6, 35] construct different views for graphs through data augmentation or prior knowledge and then maximize their agreement. It is worth noting that most graph pre-training methods are exclusively applicable to graphs with homophily, only a few works [19, 35, 34] have taken into account the extensive heterophily on graphs. Meanwhile, there is an inherent gap between the objectives of pre-training and fine-tuning phase. Pretexts extract general knowledge from graphs without supervision, while downstream tasks are rooted in specific labels. This gap may have a negative impact on the knowledge transfer [30] and hurt the performance of downstream tasks.
@@ 2.3 Graph Prompt Learning
Originated from NLP, “pre-train, prompt” paradigm [18] reformulates downstream tasks into the same template of pretexts and designs prompts for downstream adaptation. Prompt learning fully unleashes the potential of pre-trained models where adjusting only a few parameters can achieve great results with limited labels. Inspired by the success of prompt learning in nature language [1, 31] and visual data [9, 10], the graph domain has shifted the focus towards the “pre-train, prompt” paradigm [26]. GPPT [24] and GraphPrompt [20] both align the objective of downstream tasks with link prediction. ProG [25] reformulates different-level tasks to graph level and further introduces virtual prompt graphs with meta-learning. GPF [3] injects learnable prompts into the feature space for downstream task adaptation while VNT-GPPE [27] adjusts pre-trained graph transformers by prompting virtual nodes. Similar to graph pre-training, existing graph prompting methods only focus on homophilous graphs while overlook heterophilous graphs, which restricts their wide applications in the real word. Meanwhile, they heavily rely on the randomly initialized prompts, and the added virtual features or nodes are difficult to interpret. All of them ignore the rich information inherent in graph itself and the potential of prompt within models.
@@ 3 Preliminaries
3.0.2 Pre-train, Fine-tune.
The goal of graph pre-training is to learn a GNN encoder $f_{\theta}$ parameterized by $\theta$ to map each node $v$ to a low-dimensional representation $\mathbf{h}_{v}$ through the pretext in the pre-training phase, then $\mathbf{h}_{v}$ are fine-tuned according to the requirements of various downstream tasks.
3.0.3 Graph Contrastive Learning (GCL).
GCL with smoothing [36] is one GCL schema which ensures that neighboring nodes the have similar representations, the objective of its pretext is:
@@
where $\mathbf{h}_{v},\mathbf{h}_{v^{+}},\mathbf{h}_{v^{-}}$ are the representations of nodes $v,v^{+}$ and $v^{-}$ obatined by $f_{\theta}(G)$, and $\tau$ is the temperature hyper-parameter. $\mathcal{N}(v)$ is the positive sample set containing one-hop neighborhoods of node $v$ and $\mathcal{V^{-}}$ is the negative sample set randomly sampled. The illustration for this scheme of smoothing is presented in Fig. 1(a). Another GCL schema [39, 35] shown in Fig. 1(b) learns representations by contrasting graphs under augmented views. Specifically, the representation of node $v$ in one augmented view is learned to be close to that from the other view and far away from the representations of other nodes. Given two augmented $G_{1}$ and $G_{2}$ from $G$, the objective of its pretext is:
@@ 4.1 Pre-training Method
For graph pre-training, contrastive methods provide broader applicability and overlapping task sub-spaces for better knowledge transfer [21]. However, most GCL schemes explicitly or implicitly follow the homophily assumption [17, 5] and empirically face issues on heterophilous graphs. So we choose the Graph Asymmetric Contrastive Learning (GraphACL) [34] as our backbone to accommodate graph with heterophily and homophily. As shown in Fig. 1(c), GraphACL introduces the asymmetric framework with a projector. Unlike traditional symmetric schemes, the design of projector allows neighboring nodes to have different representations, and it has been theoretically and experimentally proven in [34] to capture high-order neighbor similarities on graphs and effectively handle the heterophily. The pretext loss of GraphACL is:
@@
where $\mathbf{z}_{v}=g(\mathbf{h}_{v})$ and $g(\cdot)$ can be a linear transformation layer or a MLP. The positive sample set is from $\mathcal{N}(v)$ and random data augmentation are not required. $\mathbf{h}_{v}$ and $\mathbf{h}_{v^{+}}$ are obtained separately from two encoders $f_{\theta}$ and $f_{\delta}$ updated by exponential moving average to model more complex patterns and prevent $g(\cdot)$ from degenerating to the identity function [34]. The projector enables the model to capture higher-order context information of nodes. Similar to GCL with augmented views, GraphACL also abandons the projector and solely uses the output of encoder for downstream tasks. However, the projector also contains rich pre-trained knowledge, simply discarding it would be inappropriate. It can be combined with prompt techniques to empower the downstream tasks.
@@ 4.2 Prompt Template
Prompt learning in NLP reformulates both pretext and various downstream tasks into the unified template of masked language modeling [1, 18]. As for GCL, the core is to maximize the mutual information between positive sample pairs and minimize that between negative sample pairs. To align the pretext with downstream tasks, we focus on the most common downstream tasks on graphs: link prediction and node classification, and propose a unified prompt template, denoted as $p(\cdot)$. The template transforms various tasks into pairwise similarity calculation between tokens:
@@
In the pre-training phase, the targets are set as the positive and negative samples for the anchor nodes. We use the dot product in $p(\cdot)$ for simplicity and the objective of pretext is to minimize the distance between the positive sample pairs and maximize that between the negative sample pairs in the representation space as illustrated in Eq. 3.
@@
For link prediction, the targets are set as adjacent and non-adjacent nodes for the anchor nodes. The objective is to minimize the distance between pairs of adjacent nodes and maximize that between pairs of non-adjacent nodes in the representation space.
@@
Unlike traditional node classification, we set a prototype for each class rather than predict solely by direct labels. Through class prototypes as targets, the node classification can be reformulated into pairwise similarity calculation. The objective is to minimize the distance between nodes and the intra-class prototype, while maximize that between nodes and the inter-class prototype. Formally, we denote the set of class prototype tokens as $\mathcal{C}=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{C}\}$, where $C$ is the number of classes. Each prototype token shares the same dimension with the node token and preserves class-specific semantics. Each prototype is initialized by the mean token of labeled nodes for each class [24, 20, 21]:
@@ 4.3 Self-Prompt Architecture
Most “pre-train,fine-tune” approaches for graphs introduce an additional layer for the downstream fine-tuning. Meanwhile, the projector in GCL plays a crucial role [5] in the pretext and contains partial knowledge obtained during pre-training. For example, the projector in GraphACL [34] is to preserve the context information of nodes and model local neighborhood distribution beyond homophily assumption. However, the projector is always discarded after pre-training. This motivates us to further reuse the projector as the self adapter and construct a naturally parameterized architecture for tuning. Given the input graph $\mathcal{G}$, an encoder $f_{\theta}$ parameterized by $\theta$ and the projector $g_{\phi}$ parameterized by $\phi$, the pre-training objective can be formally formulated as:
@@
where $\mathcal{L}_{dow}$ is optimized by the supervision and will be introduced in Section 4.4. Note the parameters of projector $g_{\phi}$ are initialized from pre-training stage, denoted as $g_{\phi}^{*}$. By reusing the projector as a self adapter for downstream tasks, the pre-trained knowledge can be retained and smoothly transferred to the downstream, greatly improving the efficiency of tuning for specific tasks without additional parameters.
@@ 4.4 Self-Prompt Injection and Tuning
Although the objectives of pretext and downstream tasks are aligned, task-specific prompt is still required for downstream adaptation. Currently, most works introduce trainable vectors with random initialization for prompting. The specific forms include injecting prompts in the feature space [3] and representation space [20], or using virtual nodes [27] or graphs [25] as prompts. This strategy is not stable and relies on the initialization, motivating us to design more stable prompts. Considering two key elements of the graph, namely attributes and structure, we approach the problem from these two perspectives and design prompts based on the rich information inherent in the graph itself.
@@
Instead of introducing additional parameters, the original input can be modified to provide self prompts. From structural perspective, we can add or remove edges on the adjacency matrix based on prior knowledge for prompt. There have been studies [38, 16] emphasizing the importance of multi-hop neighbors on graphs. Especially, most nodes share the same label with their two-hop neighbors on most homophilous graphs [34]. Therefore, we can simply add two-hop neighbors into the adjacency matrix as self structural prompt for node classification. Formally, we define the two-hop graph $\mathcal{G}_{2}$ from $\mathcal{G}$ and its adjacency matrix is $\mathbf{A}_{2}$, where $(\mathbf{A}_{2})_{ik}=1$ if there exists $j$ such that $e_{ij}\in\mathcal{E}$ and $e_{jk}\in\mathcal{E}$. Then we feed the $\mathcal{G}_{2}$ into our frozen pre-trained encoder $f_{\theta}$ to obtain token $\mathbf{t}_{v}$ of node $v$ for tuning:
@@
Compared with graph structure, node attributes can be even more important in certain situations, such as on heterophilous graph. But how to transform the semantic of attributes into the representation space is a challenge. Inspired by [35], we construct a unit matrix $\mathbf{I}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ to replace the adjacency matrix $\mathbf{A}$ of the original graph $\mathcal{G}$. Then we feed the prompt graph $\mathcal{G}_{I}$ into the pre-trained encoder $f_{\theta}$ to obtain semantic token $\mathbf{s}_{v}$: